OpenStudy (anonymous):

are 5,4,3 mesurmets of a right tringle

OpenStudy (anonymous):

yes.

hero (hero):

3,4,5

OpenStudy (amistre64):

yes, but not in that order lol

OpenStudy (anonymous):

\(3^2+4^2=5^2 \)

hero (hero):



OpenStudy (amistre64):

5^2 + 4^2 doesnt = 3^2 :)

OpenStudy (anonymous):

\( 5^2 + 4^2= 3^2 \) for higher values of 3

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

well thats the order there asking me in the test is a tru or false question

OpenStudy (anonymous):

then false.

OpenStudy (amistre64):

3, 4, and 5 make a triangle; and you test it by short^2 + other short^2 = long^2

OpenStudy (anonymous):

it was tru lmao

OpenStudy (amistre64):

they use that tri alot in math; just to give you a heads up

OpenStudy (anonymous):

how bout 13,12,5 is that a right tringual mesurment

OpenStudy (amistre64):

is 5^2 + 12^2 = 13^2?

OpenStudy (anonymous):

so is that tru or false

OpenStudy (amistre64):

if: 5^2 + 12^2 = 13^2 is true; then we have a right tri; if flase then we dont

OpenStudy (anonymous):

yes true cos 5^2 + 12^2 = 13^2

OpenStudy (amistre64):

the whole point of these questions is to get you to use the pythagorean thrm to check with

OpenStudy (amistre64):

if a^2 + b^2 = c^2 is true; its a right as long as a and b are shorter than c

OpenStudy (anonymous):

oooo ok so as long as a and b are shorter then c is tru

OpenStudy (amistre64):

yes, and when given the measures of the triangle; just use the shorter ones for a and b

OpenStudy (amistre64):

5 and 12 are shorter than 13; sooo 5^2 + 12^2 = 13^2, if ture is a right triangle

OpenStudy (amistre64):

since 3 and 4 are shorter than 5 if 3^2 + 4^2 = 5^2 is true; its a righty